TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8

Welcome to the home of HCC TWiki Web.

This is a web-based collaboration area for Hardware Commissioning Coordination.
All events added this month are shown below.
You can also browse by a point or day of month.


Created by Main.BorisBellesia on 2007-11-20 - 19:26

Shift HC teams and Piquet numbers - 20th Nov. 2007

PDF shifts.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x11x20x19x25/shifts.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-11-19 - 20:48

RAT CCC: Powering tests of Sector45 - 19 Nov. 2007

PDF RAT_20071119.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x11x19x20x47/RAT_20071119.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-11-14 - 18:30

RAT CCC: preparation for Powering test: Sector45 - 14th Nov.

PDF RAT_20071114.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x11x14x18x29/RAT_20071114.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-11-13 - 19:09

RAT Point4 - Preparation for Powering Sector45 - 9th Nov. 2007

PDF RAT_20071112.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x11x13x19x07/RAT_20071112.pdf

Comments:

Edit, attach file or add comment


Created by Main.IvanMoyaMartinez on 2007-11-01 - 15:14

Sector 8-1

Going-on Activities

* It has been decided to repair the flexibles in the DFBAs of Sector 8-1 before the global leak tests, allocated dates are: o DFBAP repair: Wk 44 o DFBAA repair: End of wk 44 - begining of 45

* SSS Interconnection

* Inner Triplets: o Left 1: closed, being leak tested o Right 8: closed, being leak tested

* Electrical non-conformities on instrumentation cables of the power converters will be closed by the end of the week.

* The procedure for cleaning the DFBXH is currently in engineering check. Its expected that the document will be ready for the half of the next week.

Planned activities

* PIC-PC interface test - Wk 45 * Room temperature vacuum and collimator works (wks 43-46), bake-out (wks 47 -48) * ATLAS-RP cable installation - Wks 48-49 * TS/EL: o New water-cooled cable DFBXH: Wk 45 - Saturday & Wk 46 o Replacement of water-cooled cable of RQ6 in the DFBAP - Wk 44 o Week 46: Presentation of the cables and installation of the cable supports on the DFBMs in RA83. * ElQA For the ELQA it was talked today to start the test in the L1 and in the arc at the end of the week 45 and will continue the week 46 also. There will not be interference with other groups, and the conditions for the ELQA will be granted for the date. o TP4-A: Inner triplet L1 - Week 46 o TP4-A: Inner triplet R8 - Week 46

General

* Cooling water shut-down - Wks 42-44 - All the cryogeninc activities will be stopped.

Comments:

Sector 7-8

* It will be possible that the time for the consolidation time will be longer than expected. Because of this it will be possible that the leak test will not be performed

-- Main.IvanMoyaMartinez - 05 Nov 2007

Edit, attach file or add comment

Topic revision: r42 - 2007-11-01 - JacekSzkutnik
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login